# High-frequency tails in spectral densities

### v1

@@ IV Conclusions and Implications
We show that the rate of slow nonradiative relaxation is sensitive to the functional form of the peaks used to represent spectral density in numerically exact simulations. More precisely, the high-frequency range of the spectral density is determined by the tails of the peaks. Therefore, functions whose high-frequency tails decay slowly yield faster decay of electronic excitation. This finding has several important implications for using exact quantum dynamics simulations to obtain quantitatively accurate dynamics. Firstly, the methods that truncate the spectral density at high frequency will completely miss this phenomenon. Secondly, proper care needs to be taken to discretize the high-frequency range of the spectrum for the methods that require discretization. Thirdly, we reveal that the Drude-Lorenz solvent feature sets the overall timescale for the population relaxation if the underdamped Brownian functional form is chosen to represent vibrational peaks.
We give a simple recipe to adjust for this difference in decay rates of high-frequency spectral density tails and discuss its application in analog simulation when the precise functional form of the target peaks is unattainable on the device. The adjustment yields accurate target dynamics when the timescales of dephasing and population relaxation are separated and requires no additional information beyond the spectral densities of the simulator and the target.
Acknowledgements.
@@
This material is based upon work supported by the National Science Foundation under grant No. PHY-2310657. We also wish to acknowledge helpful discussions with Chang Woo Kim, John Nichol, and Ignacio Gustin.
Appendix A 2-peak SD functions
The normalization constants for the two-peak spectral density basis functions are presented below. We show the simulations only for the Ohmic (Eq. 25) and the superohmic (Eq. 26) cases in Fig. 3.
$\Lambda_{1,k}$ $=\frac{\Lambda_{2p,k}\prod_{i=\{k_{1},k_{2}\}}\gamma_{i}(\omega_{i}^{2}+\gamma_{i}^{2})(\omega_{i}-\omega_{j}+\gamma_{i})}{\sum_{i=\{k_{1},k_{2}\},j\neq i}\gamma_{i}\left(\gamma_{i}^{2}+4\gamma_{j}^{2}+\omega_{i}^{2}\right)}$ (25)
$\Lambda_{3,k}$ $=\frac{\Lambda_{2p,k}\prod_{i=\{k_{1},k_{2}\}}\gamma_{i}}{\sum_{i=\{k_{1},k_{2}\}}\gamma_{i}}$ (26)
$\Lambda_{5,k}$ $=\frac{\Lambda_{2p,k}\prod_{i=\{k_{1},k_{2}\}}\gamma_{i}}{\sum_{i=\{k_{1},k_{2}\},j\neq i}\gamma_{i}(\omega_{j}^{2}+\gamma_{j}^{2})}$ (27)
$\Lambda_{7,k}$ $=\frac{\Lambda_{2p,k}\prod_{i=\{k_{1},k_{2}\}}\gamma_{i}}{\sum_{i=\{k_{1},k_{2}\},j\neq i}\gamma_{i}\left(4\gamma_{i}^{2}\gamma_{j}^{2}+4\gamma_{j}^{2}\omega_{j}^{2}+(\omega_{j}^{2}+\gamma_{j}^{2})^{2}\right)}$ (28)
$\Lambda_{2p,k}$ $=\frac{4(\lambda_{k_{1}}+\lambda_{k_{2}})}{\pi}\left((\omega_{k_{1}}+\omega_{k_{2}})^{2}+(\gamma_{k_{1}}-\gamma_{k_{2}})^{2}\right)\left((\omega_{k_{1}}-\omega_{k_{2}})^{2}+(\gamma_{k_{1}}+\gamma_{k_{2}})^{2}\right)$ (29)
To set the peak widths in Table 1 we work backwards from the functional form of the Ohmic 2-peak functions (Eq. (20)) to ensure that resulting spectral densities from all five basis function choices are similar with a realistic choice of widths. We pair up the eight peaks creating pairs (1,2), (3,4) etc., such that the two peaks in each pair do not have significant overlap and also have similar reorganization energy parameter ($\lambda_{k}$). We then set the peak widths of $10$ cm-1 for the four odd index peaks. The widths for the other four peaks (with even indices) are set such that the reorganization energies of each individual peak of the Ohmic 2-peak SD functions match the reorganization energy parameters $\lambda_{k}$ in Table 1.
Appendix B Derivation of tail adjustment
We derive Eqs. (23) and (24) assuming both the source and target populations follow Lindblad equation.[58] In this case the population of excited electronic state changes follows the rate equation:
$\dot{p}_{1}(t)=\gamma_{-}p_{0}(t)-\gamma_{+}p_{1}(t),$ (30)
where $\gamma_{-}$ and $\gamma_{+}$ are the (constant) rates of absorption and total (stimulated plus spontaneous) emission given in terms of spectral density as follows:
$\gamma_{-}$ $=$ $\gamma(-\Omega)=\frac{2\pi\alpha_{x}^{2}J(\Omega)}{e^{\beta\hbar\Omega}-1}=e^{-\beta\hbar\Omega}\gamma_{+}$ (31)
$\gamma_{+}$ $=$ $\gamma(+\Omega)=\frac{2\pi\alpha_{x}^{2}J(\Omega)}{1-e^{-\beta\hbar\Omega}}=2\pi\alpha_{x}^{2}J(\Omega)+\gamma_{-}$ (32)
After elimination of $p_{0}(t)$ from Eq. (30) using $p_{0}(t)+p_{1}(t)=1$ (i.e., ground and excited electronic state populations add up to 1), the equation becomes
$\dot{p}_{1}(t)=\gamma_{-}[1-p_{1}(t)]-\gamma_{+}p_{1}(t)=\gamma_{-}-\gamma p_{1}(t),$ (33)
where $\gamma=\gamma_{-}+\gamma_{+}$. Taking the following anzatz:
$p_{1}(t)=ce^{-bt}+d$ (34)
with $c$ set by initial condition as $c=p_{1}(0)-d$. Plugging it into Eq. (33):
$-bce^{-bt}=\gamma_{-}-\gamma ce^{-bt}-\gamma d,$
so $b=\gamma$ and $d=\frac{\gamma_{-}}{\gamma}=\frac{e^{-\beta\hbar\Omega}}{1+e^{-\beta\hbar\Omega}}=p_{1}(\infty)=p_{1}^{eq}$. Thus,
$p_{1}(t)=[p_{1}^{0}-p_{1}^{eq}]e^{-\gamma t}+p_{1}^{eq},$ (35)
where $p_{1}^{0}=p_{1}(0)$ is the initial population of the excited state.
To derive Eq. (23) we write Eq. (35) for both source and target, keeping in mind that initial and final excited state populations of source and target are equal, i.e. $p_{\text{target}}^{0}=p_{\text{source}}^{0}=p_{0}$ and $p_{\text{target}}^{eq}=p_{\text{source}}^{eq}=p_{eq}$
$p_{\text{target}}(t)$ $=$ $[p_{\text{target}}^{0}-p_{\text{target}}^{eq}]e^{-\gamma_{\text{target}}t}+p_{\text{target}}^{eq}$ (36)
$=$ $[p_{0}-p_{eq}]e^{-\gamma_{\text{target}}t}+p_{eq}$
$=$ $p_{\text{source}}(t)e^{-[\gamma_{\text{target}}-\gamma_{\text{source}}]t}+p_{eq}\left[1-e^{-[\gamma_{\text{target}}-\gamma_{\text{source}}]t}\right]$
$=$ $p_{eq}+\left[p_{\text{source}}(t)-p_{eq}\right]e^{-[\gamma_{\text{target}}-\gamma_{\text{source}}]t}$
Finally, Eq. (24) follows from Eq. (31) as
$\gamma=\gamma_{-}+\gamma_{+}=2\pi\alpha_{x}^{2}J(\Omega)\coth{\left(\frac{\beta\hbar\Omega}{2}\right)}$ (37)

### v2

@@ 4 Conclusions
These findings have several implications for the calculations of such tail-dominated population relaxation, for instance, when calculating nonradiative relaxation rates of realistic molecules. Firstly, the methods that truncate the spectral density at high frequencies will completely miss this phenomenon. Secondly, proper care must be taken to discretize the high-frequency range of the spectrum for the methods that require discretization. Thirdly, we reveal that the Drude-Lorentz solvent feature sets the overall timescale for the population relaxation if the underdamped Brownian functional form is chosen to represent vibrational peaks, as is commonly done.
We give a simple recipe for adjusting for this difference in the decay rates of high-frequency spectral density tails and discuss its application in analog simulation when the precise functional form of the target peaks is unattainable on the device. This adjustment expands the reach of such simulators to an arbitrary form of SD tails, provided the timescales of dephasing and population relaxation are separated. Additionally, under the same requirement of timescale separation, the adjusted results can be used to directly compare simulations performed with different functional forms of the peaks. One apt example is the adjustment of an HEOM simulation performed with the (numerically advantageous) Lorentzian cutoff to the exponential cutoff, which is challenging for HEOM calculations, but common in MCTDH and other methods.
{acknowledgement}
@@
This material is based upon work supported by the National Science Foundation under grant No. PHY-2310657. The authors thank Chang Woo Kim, John Nichol, and Ignacio Gustin for helpful discussions.
{suppinfo}
The supplementary information containing detailed descriptions of two-peak spectral density functions and the derivation of relaxation rate adjustments is available free of charge.